OpenStudy (anonymous):

solve the following system of equations. express your answer as an ordered pair in the format (a,b).

OpenStudy (anonymous):

3x+4y=16 -4x-3y=-19

OpenStudy (texaschic101):

is there a certain way you have to solve this...elimination or substitution ?

OpenStudy (anonymous):

you have to graph it and you'll end up getting (4,1)